# On complexity of alternating link equivalence

### v1

@@ 4.2. Algorithm
Step 3. Determine all inner sides of squares of $\mathcal{C}_{i},i=1,2$, and label them with $i_{1}$ for “inner”. Then label the squares that contain the already labelled squares but not other squares by $i_{2}$ on the side where the $i_{i}$ squares are. Continue until all squares are labelled this way. (Note: we might have a square labelled by $i_{1}$ on both sides, and then it is the only square in $D_{i}$.) See Figure 3 (A) for an example of such labelling.
Step 4a. When $j=i_{1}$, to compare tangles $T_{1},T_{2}$, make a pseudograph $\Gamma_{i}$ out of $T_{i},i=1,2,$ as follows. Ignore the parts of $D_{i}$ outside of $T_{i}$. In $T_{i}$, put a vertex instead of every crossing, and an edge instead of every diagram edge. For the four free ends of $T_{i}$, put a vertex of valence 1 on each (in $D_{i}$, the vertex lies on the intersection of the square encircling $T_{i}$ and the diagram edge), and give them labels $N,S,E,W$ in $T_{1}$. Then compare the resulting plane pseudographs of valence at most 4. If two pseudographs are the same, compare crossing information for the respective tangles (since the tangles are alternating, it is enough to compare just one pair of crossings, one crossing from one tangle, and one from another). If it is the same, the tangles have the same color. Label the four vertices on tangle $T_{1}$ “free ends” by $N,S,E,W$ according to the equivalence with the vertices of tangle $T_{1}$.
We now have all tangles on the $i_{1}$-sides of squares colored according to the equivalence of tangles. Make each colored tangle together with the square that encircles it into a colored graph vertex, otherwise leaving the diagram $D_{i},i=1,2$, untouched. Denote the resulting graph/diagram by $D^{\prime}_{i},i=1,2,$ respectively. The vertices of $D^{\prime}_{i},i=1,2,$ that are not colored all get the same (not previously used) color.
Step 4b. Repeat the following until we exhaust all values from $i_{2},...,i_{k}$ for $j$. Change $j$ to the next label in the list of $i_{2},....i_{k}$. By now, there are colored vertices inside any square, say $c$, that we consider. Then $c$ must encircle a product region by Lemma 1 (and all of it, since the product regions inside and outside of $c$ are not compatible by the lemma). We compare two product regions $P_{1},P_{2}$ in $D_{1}^{\prime},D_{2}^{\prime}$ respectively, as follows. Compare the list of colors of tangles inside $P_{1},P_{2}$ up to (possibly reverse) cyclic order, and compare the total number of free crossings for $P_{1},P_{2}$. If these numbers and lists are equal, also compare the lists of $N,S,W,E$ labels for free ends of equivalent tangles inside $P_{1},P_{2}$, and the crossing information for free crossings. If all coincides, the product regions are considered equivalent. Compare all pairs of product regions for the label $j$, and give the equivalent ones the same color. Make these newly colored product regions into new colored vertices, changing $D^{\prime}_{i},i=1,2$. Label the four vertices of $P_{1},P_{2}$ “free ends” by $N,S,E,W$ according to the equivalence between $P_{1}$ and $P_{2}$.
@@
Figure 3. In Figure (a), there is an example of a link diagram $D_{1}$ (in black) with a characteristic collection $\mathcal{C}_{1}$ of squares (in blue), where inner sides of squares are labeled by $i_{1},i_{2}$. Substitute 2-tangles on the sides of squares of $\mathcal{C}_{1}$ labeled by $i_{1}$ with colored graph vertices (red and green). Substitute remaining crossings of 2-tangles on the sides of squares of $\mathcal{C}_{1}$ labeled by $i_{2}$ with black vertices.
Step 5. At this step, we have no more squares. Make a colored pseudograph of what remains of $D^{\prime}_{i},i=1,2$, as in Step 4a. Compare two such vertex colored plane pseudographs.
Making pseudographs is illustrated in an example in Figure 3 (B) for $j=i_{2}$, with two blue squares encircling colored pseudographs made out of tangles.
@@ 5. Algorithmic complexity
The problem alternating link equivalence is in class P. In particular, the above algorithm has a polynomial number of steps in terms of $\text{max}\{n_{1},n_{2}\}$.
@@
The algorithm in Section 4 is for connected prime alternating reduced diagrams. We first discuss its complexity, and then its generalization to any alternating diagrams.
@@
First, detect all essential squares. For this, take all edges of $D_{i},i=1,2$. There are $2n$ of them for each diagram. Then take all (unordered) combinations of 4 edges out of them: there are (${2n\choose 4}=O(n^{4})$ of them. For each quartet of edges, check if the curve intersecting them is connected. If it is, it is a square. Check if there is more than one crossing inside and outside this square. If yes to both, then the square essential. Therefore, all essential squares can be detected in polynomial number of steps in terms of $n$, and there are at most $O(n^{4})$ of them.
@@
Second, choose characteristic squares out of the essential squares. For this, for each pair of essential squares (there are at most $O(n^{8})$ such pairs) check if they intersect as curves, and if the intersection contains a crossing. If the answer for a pair is “yes”, the squares are not characteristic. If the answers for a square have been “no” in every pair, then it is a characteristic square. All together, this takes polynomial number of steps in terms of $n$.
@@
Note that there are at most $n$ characteristic squares. Indeed, such a square $c$ must contain at least one crossing that is not contained in other characteristic squares that do not fully contain $c$.
@@
In Step 3, for each square of the characteristic collections $\mathcal{C}_{i},i=1,2$, check if there is another square fully inside it: e.g., by checking if a square shares all crossings with another square. For each of at most $4n^{2}$ pairs of sides of squares, this requires checking at most $n$ crossings, and hence yields at most $4n^{3}$ steps. This way we detect and label all $i_{1}$ sides of squares of $\mathcal{C}_{i},i=1,2$. Then for each $i_{1}$ square, consider its non-labelled side, and check whether it shares crossings with any other non-labeled side of another square. Again there are up to $4n^{2}$ pairs of sides of squares, and up to $n$ crossings for each pair to check. We have at most $4n^{3}+4n^{3}$ steps to detect $i_{1}$ and $i_{2}$ squares. Continue at most $n$ times, get at most $n$ by $4n^{3}$ steps to label all sides of squares by $i_{1},i_{2},i_{3},....i_{k},k\leq n$.
@@
Step 4 is repeated until every two tangles or product regions encircled by the squares of $\mathcal{C}$ are compared, where a tangle or a product region can lie on one of two sides of a square. So we have at most $2n$ times $2n$ pairs of tangles or product regions to compare. When we create pseudographs from fragments of a link diagram, we do this in polynomial number of steps. In the resulting pseudographs, there are no loops and there are at most two double edges between two vertices. Due to properties of a link diagram, there are at most $n$ pseudograph vertices (or $2n$ for a subdivision graph), the valence of each vertex is at most $4$, there are up to $2n$ edges for pseudograph (or $3n$ for a subdivision graph). The total number of colors is at most $n$, since we have at most $n-1$ distinct tangles, and can add one more color that is used for any pseudograph vertex that has not been obtained from a colored tangle.
@@
In Step 4b, for $j=i_{2},...,i_{k}$, the comparison for each pair of product regions is reduced to comparison of two ordered lists of colors, comparison of two numbers (i.e. the weights of product regions), and comparing $N,S,W,E$ labels for each of at most $n$ equivalent pairs of equivalent tangles. This can be done in polynomially many steps.
@@
Step 5 is polynomially reduced to vertex colored plane pseudograph isotopy with bounded color valence. It is known to have polynomial complexity as explained in subsection 2.3.
Now once there is a polynomial algorithm to compare connected prime alternating reduced diagrams, one can compare any two alternating diagrams in polynomial time. This is based on the following facts: (i) if an alternating diagram is not reduced, one can remove nugatory crossings by a 180-degree twist in ambient 3-space, and there are at most $n$ nugatory crossings; (ii) Menasco [30] proved that connected alternating diagrams represent non-split links; (iii) Menasco also proved that prime connected alternating diagrams represent prime links. Subsection 2.1 of [35] carefully describes how to use this to pass from any alternating diagrams to connected prime alternating reduced diagrams in polynomial number of steps.
6. Acknowledgements
We thank March Lackenby for helpful discussions, and Arnaud De Mesmay for helpful correspondence and encourangement. Tsvietkova is supported by NSF DMS-2005496 and NSF CAREER DMS-2142487.

### v2

@@ 5.1. Basis for the algorithm
Step 3. Determine all inner sides for squares of $\mathcal{C}_{i},i=1,2$, and label them with $\theta_{1}$. Then label the squares that contain the already labelled squares but not other squares by $\theta_{2}$ on the side where the $\theta_{1}$ squares are. Continue inductively until all sides of all squares are labelled this way by the labels from the finite list $\theta_{1},\theta_{2},\theta_{3},...$. (Note: we might have a square labelled by $\theta_{1}$ on both sides, and then it is the only square in $D_{i}$.) See Figure 4(A) for an example of such labelling.
Step 4a. Let $j=\theta_{1}$, and consider all squares labelled with $j$ at least on one side. For $j=\theta_{1}$, these are inner sides. On the inner side, there is a 2-tangle. Take a pair of such tangles $T_{1},T_{2}$ encircled by squares $c_{1},c_{2}$ in diagrams $D_{1},D_{2}$ respectively.
Comparing two 2-tangles on inner sides of squares. Make a pseudograph $\Gamma_{i}$ out of $T_{i}$ as in Remark 5. Label each free end of $T_{1}$ by N, W, S, E, starting with N on some end, and continuing counterclockwise. Then compare the resulting plane pseudographs $\Gamma_{i},i=1,2,$ up to plane isotopy. If isotopic, compare crossing information for one pair of crossings which yield equivalent vertices in $\Gamma_{1},\Gamma_{2}$, i.e. check if underpasses in $T_{1}$ correspond to underpasses in $T_{2}$. If the crossing information coincides, color the interior of a topological disk on the inner sides of $c_{1},c_{2}$ in the same color. Label the free ends of $T_{2}$ by N, W, S, E according to the isotopy with $T_{1}$.
Once we compare all squares on the inner sides as outlined in the above paragraph, we have interior of every square in $D_{1},D_{2}$ on the inner side colored. Make interior of each such square, say $c$ (together with the tangle the square encircles, say $T$) into a colored vertex $v$, otherwise leaving the diagram $D_{i},i=1,2$, untouched. Preserve the N, W, S, E labels of $T$, moving them from the ends of $T$ to the pseudograph edges incident to $v$, placing them near $v$ at those pseudograph edges (each pseudograph edge might get at most two labels this way, near each of its vertices). Denote the resulting hybrid of a colored and labelled pseudograph and a link diagram by $D^{\prime}_{i},i=1,2,$ respectively.
If at this point number of vertices of some color in $D_{1}$ is not the same as in $D_{2}$, stop with output No.
Step 4b. Repeat the following until we exhaust all values from $\theta_{2},\theta_{3},...$ for $j$.
Change $j$ to the next label in the list $\theta_{2},\theta_{3},...$. Consider a square $c_{i},i=1,2,$ in $D_{i},i=1,2,$ that has a side labelled by $j$, and refer to this side as “inside” of $c$. By now, there are subsquares inside $c_{i}$ with colored interior, and with labels N, W, S, E each.
Comparing two 2-tangles on $j$-side of squares. If there is a product region $P_{i}$ inside one of $c_{i}$ but not another, the tangles are not equivalent. If there are product regions $P_{i}$ in each of $c_{i},i=1,2$, compare the number of free vertices of $P_{i}$, and the lists of colored vertices inside $c_{i}$’s, up to cyclic order. If these numbers and lists are equal, also compare N, S, W, E labels for vertices of the same color inside $c_{i}$ as in Remark 3, and the crossing information (over/underpasses) for one pair of free crossings from $P_{1},P_{2}$. If all coincides, $c_{1}$ and $c_{2}$ are considered equivalent. Lastly, if none of the $c_{i}$ contains a product region, make colored plane pseudographs out of 2-tangles inside $c_{i}$ as in Remark 5, keeping the N, W, S, E labels as above. Compare these colored pseudographs up to plane isotopy preserving colors and labels.
Consider all squares and their sides with label $j$ in $D_{1}$ and $D_{2}$. Compare each such square in $D_{1}$ with each such square in $D_{2}$ as outlined in the above paragraph. Give the interior of equivalent squares, on the side labelled by $j$, the same color, and labels N, W, S, E for their free ends as above. This new coloring absorbs all old colors and labels from the previous steps (for previous $j$) inside the squares we consider. Make these newly colored squares into new colored vertices, keeping N, W, S, E labels for them as above. This changes $D^{\prime}_{i},i=1,2$.
If at this point the number of vertices of some color in $D_{1}$ is not the same as in $D_{2}$, stop with output No.
Step 4 is partially illustrated in Figure 4(B) for $j=i_{2}$ for one of the diagrams, say $D_{1}$.
Step 5. At this step, we have no more squares. We add an exceptional square to each of $D_{1},D_{2}$ and use the comparison procedure from Step 4b, giving Yes or No as output.
@@
Figure 4. In Figure 4(a), there is an example of a link diagram $D_{1}$ (in black) with a characteristic collection $\mathcal{C}_{1}$ of squares (in blue), where sides of squares are labeled by $\theta_{1},\theta_{2}$. Substitute 2-tangles on the sides of squares of $\mathcal{C}_{1}$ labeled by $\theta_{1}$ with colored graph vertices (red and green), as in Step 4a. We obtain $D_{1}^{\prime}$. Consider then the sides of squares labelled by $\theta_{2}$, considering them “interiors” of the squares for now. We see a product region inside the square on the left, and no product region inside the square on the right. We then turn the left square into a plane colored pseudograph. This changes $D_{1}^{\prime}$: we obtain the one in Figure 4(b). All this is a part of Step 4b.
Theorem 7.
Algorithm 6 gives $Yes$ if and only if $D_{1},D_{2}$ represent the same link, up to ambient isotopy.
Proof.
As noted above, by one of Tait flyping conjectures (proved in [19, 33, 40]), the crossing number of diagrams must be the same. This explains the stop and output No in Step 1.
Step 2 and 3 proceed without any potential stops.
By Tait flyping theorem, $D_{1}$ and $D_{2}$ are ambient isotopic if and only if they can be related by planar isotopy and a sequence of flypes.
Claim 1 We can work on the diagram in steps for different $j$, i.e. pairwise isotopy, planar or ambient rel boundary, of pairs of tangles encircled by squares of $\mathcal{C}_{1},\mathcal{C}_{2}$ (where in each pair, one tangle is from $D_{1}$, and one from $D_{2}$, and such pairs exhaust all of $D_{1},D_{2}$ and $\mathcal{C}_{1},\mathcal{C}_{2}$), is equivalent to ambient isotopy of links represented by $D_{1},D_{2}$.
Proof of Claim 1 Flypes do not move anything between interiors of different squares by Lemma 4(2), and neither does planar isotopy. Hence, $D_{1}$ and $D_{2}$ are related by a sequence of flypes and planar isotopy if and only if the tangles $D_{i}$ encircled by squares on $j$-side are related by flypes and planar isotopy, as well as are, eventually, the complements of all square interiors in $D_{i}$.
To justify steps 4-5, we proceed inductively on $j=\theta_{1},\theta_{2},...$., applying Lemma 4(3) to squares of $\mathcal{C}_{1},\mathcal{C}_{2}$ with sides labelled by the same $j$. In particular, for two such squares $c_{1},c_{2}$, let $C_{k+1}=c_{1},C^{\prime}_{k+1}=c_{2}$ in the lemma, and treat their sides labelled by $j$ as “inside”. Then let the subsquares inside them, colored on previous step (for $j-1$) be $C_{1},...,C_{k}$ and $C^{\prime}_{1},...,C^{\prime}_{k}$ respectively to apply Lemma 4 (3) statement.
By item (1) of the lemma, either $F,F^{\prime}$ do not contain products regions, or all crossings of $F$ ($F^{\prime}$) belong to one product region that is a common product region for $C_{1},...,C_{k}$ ($C_{1},...,C_{k}$). Recall that flypes can only happen in product regions.
If there are no product regions in $F,F^{\prime}$, then isotopy of plane pseudographs that respects crossing information is equivalent to planar isotopy of the respective 2-tangles by Remark 5, and to ambient isotopy rel boundary of these tangles, since no flypes are possible. If one of $F,F^{\prime}$ contains a product region, and one does not, the fragments of $D_{i}$ encircled by $c_{i}$ are not planar isotopic and the latter $c_{i}$ does not allow any flypes. Hence, by Tait flyping theorem, these cannot represent the same 2-tangles up to an ambient isotopy rel boundary. If each of $F,F^{\prime}$ contains a product region, we use Lemma 4 (1) to conclude that all of $F,F^{\prime}$ is the product region, Lemma 4 (3) to compare the tangles encircled by $c_{1},c_{2}$ and Remark 3 to compare N, W, S, E labels. Note that from working with previous $j$, we already know whether tangles in $C_{1},...,C_{k}$ and $C^{\prime}_{1},...,C^{\prime}_{k}$ are pairwise ambient isotopic rel boundary, and the isotopic ones were colored in the same color and have $N,W,S,E$ labels according to their isotopy. ∎
Remark 8.
So far we considered connected prime alternating reduced diagrams, and hence only non-split prime alternating links. Once there is a polynomial algorithm to compare non-split prime alternating links with reduced diagrams, one can compare any two alternating links given by their alternating diagrams in polynomial time. This is based on the following facts, known since 1980’s: (i) if an alternating diagram is not reduced, one can remove nugatory crossings by a 180-degree twist in ambient 3-space, and there are at most $n$ nugatory crossings; (ii) Menasco [30] proved that connected alternating diagrams represent non-split links, therefore a separating 2-sphere in a link complement corresponds to a separating curve in the link projection; (iii) Menasco also proved that connected prime alternating diagrams represent prime links, therefore the essential 2-punctured sphere in the definition of a composite link corresponds to a curve that intersects link projection twice, and has knotted arcs on both sides. Hence, there is a rapid diagram-based algorithm that makes an alternating diagram reduced, and decomposes a split composite alternating link represented by a reduced alternating diagram into connected prime parts. While [35] concerns a very different topic from what we consider here, its subsection 2.1 gives an algorithmic version of the statements about reducing diagrams, and about passing from non-prime to prime diagrams.
@@ 6. Algorithmic complexity
The problem alternating link equivalence is in class P if the input is two alternating diagrams. In particular, Algorithm 6 has a polynomial number of steps in terms of $\text{max}\{n_{1},n_{2}\}$.
@@
In the view of Remark 8, we only need to prove the second part of the theorem statement.
@@
First, detect all essential squares. For this, take all edges of $D_{i},i=1,2$. There are $2n$ of them for each diagram. Then take all (unordered) combinations of 4 edges out of them: there are (${2n\choose 4}=O(n^{4})$ of them. For each quartet of edges, check if the curve intersecting them is connected. If it is, it is a square. Check if there is more than one crossing inside and outside this square. If both hold, then the square essential. Therefore, all essential squares can be detected in polynomial number of steps in terms of $n$, and there are at most $O(n^{4})$ of them.
@@
Second, choose characteristic squares out of the essential squares. For this, for each pair of essential squares (there are at most $O(n^{8})$ such pairs) check if they intersect as curves, and if the intersection contains a crossing. If this holds for a pair, the squares are not characteristic. If this does not hold for every pair, then it is a characteristic square. All together, this takes polynomial number of steps in terms of $n$.
@@
Note that there are at most $n$ characteristic squares. Indeed, such a square $c$ must contain at least one crossing that is not contained in its proper subsquares of $\mathcal{C}_{i}$.
@@
In Step 3, for each square of the characteristic collections $\mathcal{C}_{i},i=1,2$, check if it has a subsquare: e.g., by checking if a square shares all crossings with another square. For each of at most $4n^{2}$ pairs of sides of squares, this requires checking at most $n$ crossings, and hence yields at most $4n^{3}$ steps. This way we detect and label all $i_{1}$ sides of squares of $\mathcal{C}_{i},i=1,2$. Then for each $i_{1}$ square, consider its non-labelled side, and check whether it shares crossings with any other non-labeled side of another square. Again there are up to $4n^{2}$ pairs of sides of squares, and up to $n$ crossings for each pair to check. We have at most $4n^{3}+4n^{3}$ steps to detect $i_{1}$ and $i_{2}$ squares. Continue at most $n$ times, get at most $n$ by $4n^{3}$ steps to label all sides of squares by $i_{1},i_{2},i_{3},....i_{k},k\leq n$.
@@
Step 4 is repeated until every two tangles encircled by the squares of $\mathcal{C}$ are compared, where a tangle may lie on one of two sides of a square. So we have at most $2n$ times $2n$ pairs of tangles or product regions to compare. When we create a pseudograph from a 2-tangle, we do this in polynomial number of steps. In the resulting pseudographs, there are no loops and there are at most two double edges between two vertices. Due to properties of a link diagram, there are at most $n$ pseudograph vertices (or $2n$ for a subdivision graph), the valence of each vertex is at most $4$, there are up to $2n$ edges for pseudograph (or $3n$ for a subdivision graph). The total number of colors is at most $n$, since we have at most $n-1$ distinct tangles, and can add one more color that is used for any pseudograph vertex that has not been obtained from a colored tangle.
@@
In Step 4b, for $j=i_{2},...,i_{k}$, the comparison for each pair of product regions is reduced in some cases to comparison of two ordered lists of colors, comparison of two numbers (i.e. the weights of product regions), and comparing N, W, S, E labels for each of at most $n$ equivalent pairs of equivalent tangles. This can be done in polynomially many steps. In other cases, it is reduced to vertex colored plane pseudograph isotopy with bounded color valence. It is known to have polynomial complexity as explained in subsection 3.3. Additionally, some edge labels need to be compared, which is still polynomial (N, W, E, S). Denote the larger of these polynomials by $P$.
@@
Step 5 complexity is bounded by the same polynomial $P$ as Step 4b. ∎